Overview

In this assignment, you will take on the role of a senior member of the finance team assigned to lead the investment committee of a medium-sized telecommunications equipment manufacturer. Your team is evaluating a “make-versus-buy” decision that has the potential to improve the company’s competitiveness, but which requires a significant capital investment in new equipment. The assignment is organized into two parts:

Part A: Data calculations based on the information in the scenarios Part B: Recommendations based on the calculations

Opportunity Details

The new equipment would allow your company to manufacture a critical component in-house instead of buying it from a supplier. This capability would help you stabilize your supply chain (which has suffered from some irregularities and quality issues in the past). It could also have a positive impact on profitability through the absorption of fixed costs since this new machine will have plenty of excess capacity. There may even be a possibility that the company could leverage this capability to create a new external revenue stream by providing services to other companies.

The company has been growing steadily over the past 5 years, and the financials and future prospects look good. Your CEO has asked you to run the numbers. After doing some digging into the business, you have gathered information on the following:

  • The estimated purchase price for the equipment required to move the operation in-house would be $750,000. Additional net working capital to support production (in the form of cash used in Inventory, AR net of AP) would be needed in the amount of $35,000 per year starting in year 0 and through all years of the project to support production as raw materials will be required in year o and all years to run the new equipment and produce components to replace those purchased from the vendor..

  • The current spending on this component (i.e. annual spend pool) is $1,200,000. The estimated cash flow savings of bringing the process in-house is 20% or annual savings of $240,000. This includes the additional labor and overhead costs required.

  • Finally, the equipment required is anticipated to have a somewhat short useful life, as a new wave of technology is on the horizon. Therefore, it is anticipated that the equipment will be sold after the end of the project (the last year of generated cash flow) for $50,000. (i.e. the terminal value)

  • Input from Stakeholders

    As part of your research, you have sought input from a number of stakeholders. Each has raised important points to consider in your analysis and recommendation. Some of the points and assumptions are purely financial. Others touch on additional concerns and opportunities.

QUESTION

  1. Andrew, your colleague from Accounting, recommends using the base assumptions above: 5-year project life, flat annual savings, and 10% discount rate. Andrew does not feel the equipment will have any terminal value due to advancements in technology.

Using the data presented above (and ignoring the extraneous information), for this profit and supply chain improvement project, calculate each of the following (where applicable):

  • Nominal Payback

  • Discounted Payback

  • Net Present Value

  • Internal Rate of Return

*These are the answers

Nominal Payback: 3.83

Discounted payback: 3.34

NPV: (7,888.71)

IRR: 9.61 %

Please show the steps to get to those answers. -Thanks
Answer #1

Solution:

Cash flow of 5 years

0 1 2 3 4 5
Initial Investment ( 750000)
Working Capital (35000)
Annual savings 240000 240000 240000 240000 240000
Recovery of Working Capital 35000
Total cash flow (785000) 240000 240000 240000 240000 275000
Dicounted Total cash flow (10%) (785000) 218182 198347 180316 163923 170753

1. Nominal payback

= 3+ 65000/240000 = 3.27

2. Discounted payback =

= 4+ 24232/170753 = 4+.14 =4.14

3. NPV

= 931521 - 785000 = 146521

4. IRR

IRR ia the discount rate at which NPV of the project = 0, that means present value of all inflow= present value of all outflow

so here IRR is more than 10%

= Lower rate + NPV ar Lower rate /NPV of Higher rate * ( Higher rate - Lower rate)

if 18%

NPV = 203390+ 172364+ 146071+ 123789+ 120205 =765819

= 765819- 785000 = 19181

If 16%

NPV =206897+ 178359+153758+132549+130931 = 802494

= 802494 - 785000 = 17494

IRR = 16 + 17494/19181 * 2 ( 18 - 16)

= 16+1.82 = 17.82